OpenStudy (anonymous):

Find the area of a triangle with a base length of 3 units and a height of 4 units. A.3 units B.6 units C.7 units D.8 units

OpenStudy (anonymous):

Area of the triangle = Base length x Height length / 2 A = 3x4/2 A = 12/2 A = 6 units

OpenStudy (anonymous):

thanks @GodlyMaths